how do i get the two answers for x^2 - 2x-15=0
Answer by rapaljer(4671)

This is a quadratic equation (because it has an x^2 in it!}. You have to have the equation equal to zero, which it is, then you have to factor it:

x^2 -2x-15=0
(x-5)(x+3)=0

This is what gives you TWO solutions. Set each factor equal to zero:
x-5=0
x=5

Then,
x+3=0
x=-3

For additional explanation on this topic, try my own website. Click on my tutor name "Rapaljer" anywhere in algebra.com. This takes you to my Homepage. From my Homepage, look for the link "Basic, Intermediate, and College Algebra: One Step at a Time." Choose "Basic Algebra", then look in "Chapter 2" for "Section 2.09 Solving Quadratic Equations by Factoring."

My explanations are written for students who usually have trouble understanding math, and my own students say they can understand it a LOT easier than traditional textbooks. It is also supported by my "MATH IN LIVING COLOR" pages in which I solved the hardest problems IN COLOR! If you like my website, please recommend it to family and friends. It's all FREE!!

One more thought, if you had trouble with the factoring of this problem or any other problems with factoring, please see my FACTORING sections, which are near the beginning of this very same chapter. If you need help with factoring, I even have a free video of me teaching this lesson in my own classroom a few years ago. The video is also available FREE on my website!!

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus